# Insuring uninsurable income

### v15

@@ 5 Dynamic-programming verification
$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t})}\{(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)]\},$ (10)
where $\Lambda(\lambda_{t})$ is a set of measurable controls satisfying incentive and promise-keeping, and $\lambda_{t+1}$ is given by the function describing the change of the state variable $f:\mathbb{R}_{++}\times\mathbb{R}\times\mathbb{N}\cup\{0\}\to\mathbb{R}_{++}$ such that $f(\lambda_{t},\tau(\cdot|\lambda_{t}),t)=\lambda_{t+1}$ defined in (6).
Our candidate solution is $V$ in (11) with control $\tau^{\ast}$.
@@ 6 Conclusion
This paper shows that the grim ”immiseration“ prediction of Green [1987] and Thomas and Worrall [1990] is not inevitable. A simple risk shifting rule that postpones risk can deliver near full within-period insurance while keeping promised utilities bounded. Provided the initial promise satisfies a transparent inequality, inter-generational cooperation sustains the mechanism, and every cohort attains higher lifetime utility than under autarky, with no within-cohort inequality.
Appendix A Proof of propositions 1 and 2
@@
With respect to $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ defined throughout in (5), the propositional statement is shown by Marcet and Marimon [1992]. For completeness, we report the entire proof.
@@
We proceed to show that if condition (8) holds, then $\mathcal{M}$ is Pareto optimal and not dominated by any other sequentially incentive-compatible mechanisms. By the first part of this proof, $\mathcal{M}$ is sequentially incentive compatible. By Lemma 1 since condition (8) holds, the sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$ always defined in (5). Hence $\tau^{*}$ in $\mathcal{M}$ corresponds to transfers that solve the problem (1) in every period. Therefore $\mathcal{M}$ is a Pareto optimal transfer mechanism.
It remains to prove that under condition (8) $\mathcal{M}$ is not Pareto dominated by any other sequentially incentive compatible mechanism. Suppose, contrary to our claim, that there exists a sequentially incentive compatible mechanism $\Gamma$ that Pareto dominates $\mathcal{M}$ for a given income state $e$. Let $(v_{1}^{*},v_{2}^{*})$ be the present value achieved by $\Gamma$. Set $\lambda_{0}=\overline{v}_{1}^{-1}(v_{1}^{*})$. Since condition (8) holds for any $\lambda>0$, this $\lambda_{0}$ satisfies $F(\lambda_{0})<S$ as in the proof of Lemma 1. We may now use $(\lambda_{0},e)$ as the initial condition for $\mathcal{M}$ under condition (8). Then, by construction, each individual has the same present value for both contracts. Since $\Gamma$ Pareto dominates $\mathcal{M}$, its Pareto dominance requires that $v_{2}^{*}>v_{2}(\lambda_{0},e)$. However, this contradicts the fact that solutions of $\mathcal{M}$ are Pareto optimal under the condition (8). ∎
@@ Appendix B Proof of Lemma 2
Appendix C Proof of Theorem 3
The proof of Theorem 3 uses the following lemma.
@@ Appendix C Proof of Theorem 3
Under the assumptions (i), (ii), (iii) in Theorem 3, the second-order polynomial approximation of $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is a decreasing sequence.
@@
First we consider the case where $\{\lambda_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ can be defined in (5). Using the third-order Taylor polynomial, for a given promise $\lambda_{n-1}>0$ and a given income $e\in E$, the corresponding transfer is written as
@@
It follows that the third order polynomial approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\in\mathbb{N}}$ is a decreasing sequence.
@@
We prove Theorem 3 below.
Proof of Theorem 3.
@@ Appendix E Proof of Proposition 1
$=\lambda_{t}\big{[}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}+v_{2}(\lambda_{t},e_{t})-\overline{v}_{2}(\lambda_{t})$
and we see that the last terms are maximized by $\tau^{*}$. It follows that (E.1) is written as
$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t})}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)],$
and we see the desired result.

### v16

@@ 5 Dynamic-programming verification
For a given current promise $\lambda$, let $\Lambda(\lambda)$ denote the set of measurable controls $\tau:E\to\mathbb{R}$ such that the implied continuation promise $\lambda^{\prime}_{\tau}(e\mid\lambda)$ defined by (5) whenever feasible, and by (6) otherwise, is well defined for all $e\in E$, and the pair $(\tau,\lambda^{\prime}_{\tau})$ satisfies the promise-keeping condition (3) and the sequential incentive constraint (4).
@@
$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t})}\{(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda^{\prime}_{\tau}(e_{t}|\lambda_{t}),t+1)]\}.$ (10)
Under Assumption 1, $\lambda^{\prime}_{\tau}$ is given by (5).
Our candidate solution is $V$ in (11) with control $\tau^{\ast}$. [^4] Although we write $V(\lambda_{t},t)$ to match the terminal-condition formulation of Wiszniewska-Matyszkiel [2011], the economic environment is stationary; the time argument is used only for the dynamic-programming verification. Along the candidate mechanism, $V(\lambda_{t},t)=v_{1}(\lambda_{t})$.
@@ 6 Conclusion
This paper shows that immiseration is not unavoidable within the risk-shifting mechanism studied here. The mechanism induces a bounded promised-utility process and, under the paper’s sufficient condition, is sequentially efficient. In the stationary OLG extension, the paper also identifies sufficient conditions under which intergenerational transfers sustain the mechanism and can improve on autarky.
Appendix A Proof of theorems 1 and 2
@@
Similar proof can be found in Marcet and Marimon [1992].
@@
We next show that, under condition (8), the continuation rule induced by $\mathcal{M}$ solves problem (1) at every feasible current state. By the first part of this proof, $\mathcal{M}$ is sequentially incentive compatible. By Lemma 1 since condition (8) holds, the sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$ always defined by (5). Hence $\tau^{*}$ in $\mathcal{M}$ corresponds to transfers that solve the problem (1) in every period. Therefore, conditional on the current promised state and within the admissible class where (5) applies, $\tau^{*}$ solves the planner’s static problem at each date and induces a Pareto-efficient continuation payoff pair.
It remains to prove that under condition (8) $\mathcal{M}$ is not Pareto dominated by any other sequentially incentive compatible mechanism. Suppose, contrary to our claim, that under Assumption 1 there exists a sequentially incentive compatible mechanism $\Gamma$ that Pareto dominates $\mathcal{M}$ for a given income state $e$. Let $(v_{1}^{*},v_{2}^{*})$ denote the continuation payoff pair induced by $\Gamma$. Since $v_{1}^{*}$ lies in the range of $\overline{v}_{1}$, define
$\lambda_{0}=\overline{v}_{1}^{-1}(v_{1}^{*}).$
Because condition (8) depends only on the primitives and holds for every $\lambda>0$, it also holds at $\lambda_{0}$. Hence $\mathcal{M}$ is well defined from the initial state $(\lambda_{0},e)$, and its continuation rule is given by (5). By construction, $\Gamma$ and $\mathcal{M}$ deliver the same continuation utility to the agent, so $v_{1}^{*}=\overline{v}_{1}(\lambda_{0})$. Since $\Gamma$ Pareto dominates $\mathcal{M}$, it must satisfy
$v_{2}^{*}>v_{2}(\lambda_{0},e).$
Hence
$\lambda_{0}v_{1}^{*}+v_{2}^{*}>\lambda_{0}\overline{v}_{1}(\lambda_{0})+v_{2}(\lambda_{0},e),$
which contradicts the optimality of $\tau^{*}$ in problem (1) at $(\lambda_{0},e)$. ∎
@@ Appendix B Proof of Lemma 2
Appendix C Proof of Proposition 1
The proof of Proposition 1 uses the following lemma.
@@ Appendix C Proof of Proposition 1
Under the assumptions (i), (ii), (iii) in Proposition 1, the second-order polynomial approximation of $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is a decreasing sequence.
@@
First we consider the case where $\{\lambda_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ can be defined in (5). We use a third-order Taylor expansion to obtain the second-order approximation used in the proof of Proposition 1. Using the third-order Taylor polynomial, for a given promise $\lambda_{n-1}>0$ and a given income $e\in E$, the corresponding transfer is written as
@@
It follows that the second-order polynomial approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\in\mathbb{N}}$ obtained from the third-order expansion is a decreasing sequence.
@@
We prove Proposition 1 below.
Proof of Proposition 1.
@@ Appendix E Proof of Proposition 2
$=\lambda_{t}(1-\beta)u(e_{t}+\tau(e_{t}\mid\lambda_{t}))+v_{2}(\lambda_{t},e_{t})+\lambda_{t}\beta\overline{v}_{1}(\lambda_{t})-\overline{v}_{2}(\lambda_{t}).$
Thus, up to the additive constant $\lambda_{t}\beta\overline{v}_{1}(\lambda_{t})-\overline{v}_{2}(\lambda_{t})$, the Bellman integrand is exactly the objective of problem (1). Hence $\tau^{*}$ attains the supremum. Therefore $\tau^{*}$ maximizes (E.2).